OpenStudy (ammarah):

What does not represent the side length of a triangle? a. 3,3,5 b. 5,12,13 c. 8,8,8 d. 5,5,10

OpenStudy (ammarah):

@Frostbite

OpenStudy (anonymous):

add all and look for that who is not smaller than one side

OpenStudy (anonymous):

for example: Adding the 1st and 2nd, we get 3+3 = 6 and that is greater than the 3rd side 4. Adding the 1st and 3rd, we get 3+4 = 7 and that is greater than the 2nd side 3. Adding the 2nd and 3rd, we get 3+4 = 7 and that is greater than the 1st side 3. Therefore (3,3,4) can represent the sides of a triangle

OpenStudy (anonymous):

do u understand?

OpenStudy (ammarah):

wait so whats the answer to my question? is it a.?

OpenStudy (ammarah):

and the question is asking for the one that doesnt match...

OpenStudy (anonymous):

but u have to do all so and to find which is that side that is not bigger then the other side 4+5>10 so this is an example which doesn't match.

OpenStudy (anonymous):

okay lets do one of them pick on

OpenStudy (ammarah):

a.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

 this is